MODS1 Dark Commissioning and Spring SV Run: UTC 2011 May 30

Observer(s): R. Pogge (OSU), M. Wagner (OSU/LBTO), M. Pedani (LBTO)
Telescope Operator: G. Bechetti (LBTO)
Telescope Support: M. Midkiff (LBTO)
Instrument Support: D. Officer (LBTO)

Summary

No observing - high winds all night, and high humidity after about 0200 MST. Rotator appears to be functioning again. Worked through the comparison and flat-field calibration scripts, made a library of flats and mask images. Exercised observing planning tools accompanied by the sounds of winds up into the 30-40m/s range roaring outside.

Preparations

As of 1330 MST, the LDG rotator was still not functioning (see IssueTrak #3411), despite replacing the driver module during the morning. Dave Ashby and Dan Cox worked on the problem remotely from Tucson with assistance from Mike and Doug on-site.

At 1530 MST, we were informed that the rotator has been operating without problems for the past 2 hours. Doug is locking out the rotator to fill MODS1, then they will go back to testing. Of course, all afternoon the winds have been howling at >30m/s, the biggest gust we saw was ~42m/s (~94 mph). Dave Ashby says the problem was intermittent, so it is not going to be clear if the problem is solved until we try to use it a lot.

Exercised the observing preparation scripts, identifying and addressing bugs, annoyances, and confusing bits. These will be part of the modsTools suite of programs we'll release with how-to documents to the partners this summer. All of the spring SV program scripts were generated using these tools. The modsDisp FITS image autodisplay and archive disk monitor program works well. We setup the LBTO account to use the MODS observing tools transparently, following the successful use of the OSURC account to observe the past two runs (we have not operated MODS1 for daytime calibrations or night-time observing from the instrument "technical" workstations since February except for daytime testing and bug fixes).

Details

All times are UTC unless otherwise noted.

0005 - Geno firing up the LBT and doing his closed-dome checkout. Winds are still blowing strong outside, sustained 30m/s with gusts up to 37m/s during the past hour. It is mostly clear if a little hazy as the winds in the valleys below have been kicking up dust. It does not look like we'll be opening right away tonight.

0230 - Sunset, some clouds to the south mostly going away. Winds still above limits, but less intense than earlier and showing a slow diminishing trend.

0445 - Wind still above limits, but slowly trending downward. We might be able to open after midnight at this rate.

0600 - Doing some internal calibration observations while the wind is still howling.

0710 - Testing different types of spectral calibration flats. Humidity has risen to 70% in the past hour or so, and the wind is still sustained above 20m/s with occasional building-rattling gusts.

Having good success obtaining pixel-to-pixel grating flats in slitless mode with the VFLAT lamp (QTH too bright). VFLAT 5 Red+ND1.5 and Blue+Clear, short (1-5 sec) exposures works well (1.5s red, 5s blue both give ~40Kadu peak). Blue+UG5 at 2s and VFLAT 10 gives ~36Kadu peak. A particular advantage in both channels is the slitless mode doesn't have struts from the segmented long-slits in the way, so it makes for a good general solution for long-slit pixel-to-pixel zero-color flat fields for the grating. Prisms don't disperse enough for this, so they remain a challenge.

0800 - MOS mask spectral flats. 1 each of R+B QTH, R+B(UG5) QTH, and Red VFlat. Humidity now pushing up to 80%, wind trending back up into the mid to upper 20m/s range, and more building-rattling gusts than the previous hour. All the MOS masks are clean and sharp.

0900 - Completed sequence of long-slit grating flags, 3 per mask same sequence and exptimes as the MOS masks. Meanwhile humidity has been 82%, and wind gusts as high as 32m/s (~72mph). The new facility long-slit masks are very clean, URIC did a terrific job.

0930 - Humidity about 87%, winds still around 25m/s with occasional higher gusts. Nice view of the Milky Way on the all-sky camera. Finished the sequences of calibration images. Mark and Marco are heading down tomorrow and called it a night.

1000 - Called it a night 30m before the start of astronomical twilight, winds still howling, but humidity appears to have topped out and is starting to decline. 2 down, 2 to go.

-- RichardPogge - 29 May 2011
Topic revision: r5 - 30 May 2011, RichardPogge
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback